34. Starting from rest, a 25 g bouncy ball falls from a height of 10. m onto a flat, horizontal surface and rebounds with 99% of its incident speed. Find the amount of time that the ball was in contact with the ground if the average net force on the ball during the collision is 45 N.
Answer #1

ball speed just before hitting the ground

v^2=u^2+2*h*h

v^2= 0 +2*9.8*10

v =sqrt(2*9.8*10) = 14 m/s

rebound speed vf = 0.99*14 =13.86 m/s

average force = dP/dt = change in momentum/time = m( vf-vi) /t

45 = 25*10^-3 *(13.86 -( -14))/t

from here t = 0.0154778 sec =15.48 ms